OpenStudy (marcelie):

can someone tell me how i can study for series and sequences for calc2 since there are many formulas to know...

Nnesha (nnesha):

practice practice practice

OpenStudy (marcelie):

ive been reveiwing what each formula is used for but somehow i get them confuse :/

Nnesha (nnesha):

\[practice -> \infty \]

OpenStudy (marcelie):

lool okie ill do that for sure ;p

Nnesha (nnesha):

what do you mean by *review*?

OpenStudy (marcelie):

oh like reading what each series is used for

Nnesha (nnesha):

don't do that

Nnesha (nnesha):

try to solve 20+++ questions then you don't need to memorize anything :D

OpenStudy (marcelie):

ohh okay ill do that lol

OpenStudy (vuriffy):

Practice makes greatness.

OpenStudy (osprey):

I don't know what calc2 is, but below is sort of a "beginners' guide". 1+2+3 = 6 This is an ARITHMETIC progression. to get the next number you add 1 to the previous one. 1, 3, 5 add 2 to the previous number. etc. series and sequences make me go cross eyed, actually.

Nnesha (nnesha):

Quiztime!! ?? o.-

Nnesha (nnesha):

\[\sum_{1}^{\infty} \frac{ (-1)^{n+1} }{ \sqrt[n]{n} }\]

Nnesha (nnesha):

n=1 ** :=))

zepdrix (zepdrix):

Remember back earlier in the semester when you were doing weird integrals? You often had the option to do either `Integration by Parts` or `Trig Substitution`. You had multiple ways to attack one problem. But it usually turned out that one method was easier than the other. Same with series. You'll have a lot of different tools/tests that you can try. Getting familiar with which one works best in each situation just takes... practice :p Sorry I know I sound like a broken record at this point lol.

OpenStudy (vuriffy):

It's okay, I say that about every math puzzle. @zepdrix (:

OpenStudy (osprey):

i^0 =1, i^1 =i, i^2 = -1 ... i^4 =1 type territory ? i^8 =1 sort of thing ...

OpenStudy (marcelie):

thanks for the feedback :)